OpenStudy (anonymous):

Solve for x: −3|x + 7| = −12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (anonymous):

@phi

OpenStudy (texaschic101):

-3 | x + 7 | = -12 to start, we want the absolute value sign and everything in it on one side, and everything else on the other side. So we will first divide both sides by -3. giving us : | x + 7| = 4 you following me so far ?

OpenStudy (texaschic101):

| x + 7 | = 4 now we get 2 equations from this... x + 7 = 4 and -(x + 7) = 4 now lets solve them... x + 7 = 4 -(x + 7) = 4 x = 4 - 7 -x - 7 = 4 x = -3 -x = 4 + 7 -x = 11 x = -11 so our solutions are : x = -3 and x = -11

OpenStudy (anonymous):

Divide? Okay, yes I am following.

OpenStudy (texaschic101):

yes...you have to get that -3 to the other side...and to do that, we divide...thus getting rid of the -3 on the left side. leaving you with : | x + 7| = 4 then turn it into 2 equations...x + 7 = 4 and - (x = 7) = 4 and the solutions are : x = -3 and -11 any questions ?

OpenStudy (texaschic101):

oops...-(x + 7) = 4....sorry about that typo

OpenStudy (anonymous):

Okay, so our two possible solutions can be + or -?